# FuXi-Ocean: A Global Ocean Forecasting System with Sub-Daily Resolution

### v1

@@ 1 Introduction
Recently, deep learning approaches emerge as promising alternatives to traditional OGCMs. Studies [45, 43, 46, 2, 11] demonstrate that such models can achieve comparable or even superior accuracy at daily temporal resolution compared to operational numerical systems, and their operational efficiency can be improved by a thousand times. However, extending these models to sub-daily resolutions deep learning unique challenges. Ocean variables exhibit diverse temporal dynamics that vary significantly across depths and spatial regions. Surface variables, such as sea surface temperature, often follow pronounced diurnal cycles similar to atmospheric variables, whereas deeper ocean currents evolve on considerably slower timescales. Unlike atmospheric forecasting, which is dominated by high-frequency variability, ocean prediction must accommodate a broad spectrum of temporal behaviors, from fast-changing surface processes to slowly evolving deep-ocean processes. A key limitation of existing deep learning-based ocean forecasting models lies in their predominant focus on daily forecasts, which restricts their ability to adaptively model variable-specific temporal dynamics at sub-daily intervals. Lacking mechanisms to adaptively process temporal information across multiple timescales, these models often struggle with the complexity of high-frequency sub-daily predictions. Moreover, designing effective sub-daily models requires architectures capable of learning efficiently from limited historical data while avoiding overfitting to spurious correlations.
@@
In this work, we present FuXi-Ocean, the first deep learning-based global ocean forecasting model to achieve six-hour temporal resolution at eddy-resolving 1/12° spatial resolution. Our approach explicitly addresses the challenges of sub-daily ocean prediction through three key innovations: First, we design an autoregressive architecture specifically tailored to capture multi-scale temporal dependencies of different oceanic variables across an unprecedented depth range (0-1500 m). Unlike preivous models that treat all variables uniformly, our model adaptively learns temporal context appropriate for each variable and region, effectively distinguishing between fast-evolving surface processes (e.g., diurnal warming) and slowly varying deep-ocean dynamics. Second, we introduce the Mixture-of-Time (MoT) module that adaptively integrates predictions from multiple temporal windows based on their empirical reliability for each variable. This mechanism enables the model to select the most informative temporal context, thereby mitigating the accumulation of forecast errors typically encountered in sequential prediction tasks. Third, we demonstrate remarkable data efficiency, achieving state-of-the-art performance with only 9 years of training data—significantly less than required by comparable models. This efficiency stems from our architecture’s ability to effectively leverage physical constraints and spatial coherence in the learning process. Our key contributions are as follows:
@@ 3.1 Problem Formulation
We focus on five fundamental ocean state variables that collectively characterize the physical state of the global ocean: temperature (T), salinity (S), zonal and meridional components of ocean currents (U and V), and sea surface height (SSH). While SSH is inherently a surface variable, the remaining four variables exhibit substantial vertical variability that necessitates prediction across multiple depth levels. We discretize the water column into 20 strategically selected depth levels from the surface (0m) to the deep ocean (1500 m): 0, 2, 4, 6, 10, 20, 30, 40, 50, 60, 70, 80, 100, 125, 150, 200, 300, 500, 1000, and 1500 m. This vertical resolution is particularly dense near the surface to accurately capture upper-ocean dynamics that strongly influence air-sea interactions. Consequently, our input and output tensors comprise $C=81$ channels (4 variables $\times$ 20 depth levels + SSH).
@@ 3.2.1 Feature Extraction
where W represents the learnable convolution parameters and $\odot$ denotes element-wise multiplication. This modulation enhances the encoder’s sensitivity to region-specific patterns, such as boundary currents or seasonal mixed layer dynamics. The encoder processes each input time step $t-i$ ($i\in\{0,1,2,3\}$) to produce feature tensors $\text{F}^{t-i}\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$. Here, $C^{\prime}$ is the feature dimension while $H^{\prime}$ and $W^{\prime}$ represent the downsampled spatial dimensions.
@@ 3.2.3 Mixture-of-Time Module
A key innovation in our framework is the Mixture-of-Time (MoT) module, illustrated in Figure 1(c), that adaptively integrates predictions from multiple temporal windows. This module addresses a fundamental challenge in ocean forecasting—different oceanic variables exhibit distinct temporal evolution characteristics, from fast-changing surface processes to slow-evolving deep-water phenomena. The MoT module performs channel-wise adaptive selection across predictions derived from different temporal windows. Specifically, we maintain a selection matrix $V\in\mathbb{R}^{C\times 4}$ that quantifies prediction reliability across temporal scales for each channel (variable and depth combination). For a given channel $c$ and spatial location $(h,w)$, the final prediction is computed as:
@@
Here, the TopK operator selects the $K$ most reliable temporal windows (corresponding to smallest values in V) for each channel. Specifically, for each channel $c$, it assigns a value of 1 to the $K$ smallest elements in the vector $\text{V}(c,\cdot)$ and 0 to all others. $\overline{\text{X}}^{t+1}_{t-i}$ represents the forecast for time $t+1$ generated using historical context beginning at time $t-i$. This approach enables variable-specific temporal context selection without increasing model complexity. During training, we update the selection matrix based on prediction performance:
@@
where $\alpha$ controls update momentum, and MAE measures prediction error. The spatially averaged MAE captures the reliability of each temporal window for each variable, which is then normalized through softmax to ensure the selection weights sum to one. This adaptive approach enables FuXi-Ocean to dynamically adjust its reliance on different temporal windows for each variable and region. For example, deep ocean current predictions might benefit from recent observations that capture immediate flow evolution, while surface temperature predictions might rely more on longer temporal contexts that effectively represent diurnal cycles and day-night transitions.
@@ 4.1 Data
IV-TT Framework. For evaluation against real-world observations, we employ the GODAE Ocean View Intercomparison and Validation Task Team (IV-TT) Class 4 framework [39]. Details are in the appendix. This framework provides observational datasets from drifting buoys for 2022, alongside interpolated outputs from operational numerical forecasting systems. However, after evaluation, the raw data errors for salinity and temperature were too large to be used directly. We primarily evaluate sea surface temperature (SST) forecasts, as this variable exhibits high variability and effectively demonstrates forecasting skill. To ensure consistent comparison, we transform the irregularly distributed observational data to a regular grid matching FuXi-Ocean’s output format using nearest-neighbor interpolation. For operational evaluation, FuXi-Ocean is initialized with six-hourly analysis fields from HYCOM.
Figure 2: Global RMSE distribution of sea surface. From top to bottom, the results represent salinity, temperature, and ocean current UV components, and from left to right correspond to different forecast lead times. Each subplot represents the average RMSE (lower is better) for the test set.
4.2 Implement Details
@@ 6 Conclusion
We present FuXi-Ocean, the first data-driven global ocean forecasting system achieving 6-hour temporal resolution at a 1/12° spatial scale with coverage from surface to 1500 m depth. Our Mixture-of-Time module adaptively captures different temporal dependencies across ocean variables, addressing a key limitation of previous models. FuXi-Ocean outperforms operational numerical systems in predicting sea surface temperature while requiring significantly fewer computational resources and relying only ocean variables as input. Its 6-hour forecasts demonstrate superior accuracy than daily-averaged predictions from state-of-the-art numerical models, confirming the effectiveness of our approach for high-frequency ocean prediction.
@@
Appendix A Limitations
Despite these promising results, our approach has several limitations that warrant further investigation. The reliance on reanalysis data for training, while necessary given data availability constraints, introduces potential biases from the underlying numerical models. This challenge is compounded by the scarcity of publicly available high-quality sub-daily ocean reanalysis data. In future work, we plan to address this limitation by integrating HYCOM analysis fields with Argo observational data and other in-situ measurements to create a more robust training dataset less dependent on any single numerical model’s biases. Additionally, our evaluation thus far focus primarily on standard RMSE metrics, which, while useful for comparing forecast systems, may not fully capture the model’s ability to represent specific ocean phenomena like mesoscale eddies, western boundary currents, or seasonal thermocline transitions. Future analysis will incorporate phenomenon-specific evaluation metrics and physical consistency measures to provide deeper insights for model improvement. Finally, while our system’s forecast accuracy remains robust through 10 days, performance for longer-term predictions (seasonal to interannual) remains unexplored, representing another important direction for future research.
Appendix B Borader Impacts
Appendix C Additional results
C.1 Performance Test Supplement
C.2 Performance Test Supplement
Appendix D Data
Table 1: Comparison of Training Data for Existing Ocean Models
Name of the Model Training Data Time Range Input Variables
Xihe 1993-2020 SST, U10, V10, T, S, U, V, SSH
Langya 1993-2021 SST, U10, V10, T, S, U, V, SSH
WenHai 1993-2020 SST, U10, V10, T, S, U, V, SSH
FuXi-Ocean 2006-2015 T, S, U, V, SSH
Table 2: The source of observation data
Variable Data type Source of data
S In situ Argo profiles Argo GDAC (as above)
Figure 10: Global distribution map of ARGO and other buoys.
Appendix E Supplementary details

### v2

@@ 1 Introduction
Recently, deep learning approaches emerge as promising alternatives to traditional OGCMs. Studies [45, 43, 46, 2, 11] demonstrate that such models can achieve comparable or even superior accuracy at daily temporal resolution compared to operational numerical systems, and their operational efficiency can be improved by a thousand times. However, extending these models to sub-daily resolutions presents unique challenges. Ocean variables exhibit diverse temporal dynamics that vary significantly across depths and spatial regions. Surface variables, such as sea surface temperature, often follow pronounced diurnal cycles similar to atmospheric variables, whereas deeper ocean currents evolve on considerably slower timescales. Unlike atmospheric forecasting, which is dominated by high-frequency variability, ocean prediction must accommodate a broad spectrum of temporal behaviors, from fast-changing surface processes to slowly evolving deep-ocean processes. A key limitation of existing deep learning-based ocean forecasting models lies in their predominant focus on daily forecasts, which restricts their ability to adaptively model variable-specific temporal dynamics at sub-daily intervals. Lacking mechanisms to adaptively process temporal information across multiple timescales, these models often struggle with the complexity of high-frequency sub-daily predictions. Moreover, designing effective sub-daily models requires architectures capable of learning efficiently from limited historical data while avoiding overfitting to spurious correlations.
@@
In this work, we present FuXi-Ocean, the first deep learning-based global ocean forecasting model to achieve six-hour temporal resolution at eddy-resolving 1/12° spatial resolution. Our approach explicitly addresses the challenges of sub-daily ocean prediction through three key innovations: First, we design an autoregressive architecture specifically tailored to capture multi-scale temporal dependencies of different oceanic variables across an unprecedented depth range (0-1500 m). Unlike previous models that treat all variables uniformly, our model adaptively learns temporal context appropriate for each variable and region, effectively distinguishing between fast-evolving surface processes (e.g., diurnal warming) and slowly varying deep-ocean dynamics. Second, we introduce the Mixture-of-Time (MoT) module that adaptively integrates predictions from multiple temporal windows based on their empirical reliability for each variable. This mechanism enables the model to select the most informative temporal context, thereby mitigating the accumulation of forecast errors typically encountered in sequential prediction tasks. Third, we demonstrate remarkable data efficiency, achieving state-of-the-art performance with only 9 years of training data—significantly less than required by comparable models. This efficiency stems from our architecture’s ability to effectively leverage physical constraints and spatial coherence in the learning process. Our key contributions are as follows:
@@ 3.1 Problem Formulation
We focus on five fundamental ocean state variables that collectively characterize the physical state of the global ocean: temperature (T), salinity (S), zonal and meridional components of ocean currents (U and V), and sea surface height (SSH). While SSH is inherently a surface variable, the remaining four variables exhibit substantial vertical variability that necessitates prediction across multiple depth levels. We discretize the water column into 20 strategically selected depth levels from the surface (0m) to the deeper ocean (1500 m): 0, 2, 4, 6, 10, 20, 30, 40, 50, 60, 70, 80, 100, 125, 150, 200, 300, 500, 1000, and 1500 m. This vertical resolution is particularly dense near the surface to accurately capture upper-ocean dynamics that strongly influence air-sea interactions. Consequently, our input and output tensors comprise $C=81$ channels (4 variables $\times$ 20 depth levels + SSH).
@@ 3.2.1 Feature Extraction
where W represents the learnable convolution parameters and $\odot$ denotes element-wise multiplication. This modulation enhances the encoder’s sensitivity to region-specific patterns, such as boundary currents or seasonal mixed layer dynamics. The encoder processes each input time step tit-i ($i\in\{0,1,2,3\}$) to produce feature tensors $\text{F}^{t-i}\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$. Here, $C^{\prime}$ is the feature dimension while $H^{\prime}$ and $W^{\prime}$ represent the downsampled spatial dimensions.
@@ 3.2.3 Mixture-of-Time Module
A key innovation in our framework is the Mixture-of-Time (MoT) module, illustrated in Figure 1(c), that adaptively integrates predictions from multiple temporal windows. This module addresses a fundamental challenge in ocean forecasting—different oceanic variables exhibit distinct temporal evolution characteristics, from fast-changing surface processes to the relatively slow and stable subsurface variability extending down to the twilight zone ( 1500 m). The MoT module performs channel-wise adaptive selection across predictions derived from different temporal windows. Specifically, we maintain a selection matrix $V\in\mathbb{R}^{C\times 4}$ that quantifies prediction reliability across temporal scales for each channel (variable and depth combination). For a given channel $c$ and spatial location $(h,w)$, the final prediction is computed as:
@@
Here, the TopK operator selects the $K$ most reliable temporal windows (corresponding to smallest values in V) for each channel. Specifically, for each channel $c$, it assigns a value of 1 to the $K$ smallest elements in the vector $\text{V}(c,\cdot)$ and 0 to all others. $\overline{\text{X}}^{t+1}_{t-i}$ represents the forecast for time $t+1$ generated using historical context beginning at time tit-i. This approach enables variable-specific temporal context selection without increasing model complexity. During training, we update the selection matrix based on prediction performance:
@@
where $\alpha$ controls update momentum, and MAE measures prediction error. The spatially averaged MAE captures the reliability of each temporal window for each variable, which is then normalized through softmax to ensure the selection weights sum to one. This adaptive approach enables FuXi-Ocean to dynamically adjust its reliance on different temporal windows for each variable and region. For example, deeper ocean current predictions might benefit from recent observations that capture immediate flow evolution, while surface temperature predictions might rely more on longer temporal contexts that effectively represent diurnal cycles and day-night transitions.
@@ 4.1 Data
IV-TT Framework. For evaluation against real-world observations, we employ the GODAE Ocean View Intercomparison and Validation Task Team (IV-TT) Class 4 framework [39], which is obtained from publicly available sources. Details are in the appendix. This framework provides observational datasets from drifting buoys for 2022, alongside interpolated outputs from operational numerical forecasting systems. In addition, we apply a filter to remove anomalous points. Specifically, we compare HYCOM reanalysis data with observational data and remove outliers with high MAE from the observations to ensure fairness in the comparison process. However, after evaluation, the raw data errors for salinity and temperature were too large to be used directly. We primarily evaluate sea surface temperature (SST) forecasts, as this variable exhibits high variability and effectively demonstrates forecasting skill. To ensure consistent comparison, we transform the irregularly distributed observational data to a regular grid matching FuXi-Ocean’s output format using nearest-neighbor interpolation. For operational evaluation, FuXi-Ocean is initialized with six-hourly analysis fields from HYCOM.
Figure 2: Global RMSE distribution of sea surface. From top to bottom, the results represent salinity (psu), temperature (°C), and ocean current U/V components (m/s), and from left to right correspond to different forecast lead times. Each subplot represents the average RMSE (lower is better) for the test set.
4.2 Implementation Details
@@ 6 Conclusion
We present FuXi-Ocean, the first data-driven global ocean forecasting system achieving 6-hour temporal resolution at a 1/12° spatial scale with coverage from surface to 1500 m depth. Our Mixture-of-Time module adaptively captures different temporal dependencies across ocean variables, addressing a key limitation of previous models. FuXi-Ocean outperforms operational numerical systems in predicting sea surface temperature while requiring significantly fewer computational resources and relying only on ocean variables as input. Its 6-hour forecasts demonstrate superior accuracy than daily-averaged predictions from state-of-the-art numerical models, confirming the effectiveness of our approach for high-frequency ocean prediction.
@@ 80. Guidelines:
Appendix A Limitations
Despite these promising results, our approach has several limitations that warrant further investigation. The reliance on reanalysis data for training, while necessary given data availability constraints, introduces potential biases from the underlying numerical models. This challenge is compounded by the scarcity of publicly available high-quality sub-daily ocean reanalysis data. In future work, we plan to address this limitation by integrating HYCOM analysis fields with Argo observational data and other in-situ measurements to create a more robust dataset. For subsurface validation, this will involve a rigorous data matching process, where observational profiles are spatio-temporally collocated with the model grid points, and quality control filters are applied to remove outliers and inconsistent measurements before comparison. The filters use HYCOM analysis field data as an anchor to remove outlier data with high variability. Additionally, our evaluation thus far focus primarily on standard RMSE metrics, which, while useful for comparing forecast systems, may not fully capture the model’s ability to represent specific ocean phenomena like mesoscale eddies, western boundary currents, or seasonal thermocline transitions. Future analysis will incorporate phenomenon-specific evaluation metrics and physical consistency measures to provide deeper insights for model improvement. Finally, while our system’s forecast accuracy remains robust through 10 days, performance for longer-term predictions (seasonal to interannual) remains unexplored, representing another important direction for future research.
Appendix B Broader Impacts
Appendix C Additional results
C.1 Performance Test Supplement
C.2 Performance Test Supplement
Appendix D Data
Table 1: Comparison of Training Data for Existing Ocean Models
Name of the Model Training Data Time Range Input Variables
Xihe 1993-2020 SST, U10, V10, T, S, U, V, SSH
Langya 1993-2021 SST, U10, V10, T, S, U, V, SSH
WenHai 1993-2020 SST, U10, V10, T, S, U, V, SSH
FuXi-Ocean 2006-2015 T, S, U, V, SSH
Table 2: The source of observation data
Variable Data type Source of data
S In situ Argo profiles Argo GDAC (as above)
Figure 10: Global distribution map of ARGO and other buoys.
Appendix E Supplementary details